OpenStudy (anonymous):

Using the discriminant formula how will you determine the nature of the roots for x^2-4x-5 ?

OpenStudy (anonymous):

\[D = b^2 - 4ac\] a = 1, b = -4 and c= -5 \[D = 16 - 20 = -4\] Note these points: \[\large D > 0\] There will be two real and distinct roots.. \[\large D = 0\] There will be two real but equal roots. \[\large D < 0\] There will be two complex roots..

OpenStudy (anonymous):

Here D is less than 0, so which condition applies here??

OpenStudy (anonymous):

there are 2 complex roots.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Welcome dear..